# Yuqin Dai1,3,    Guoqing Wang3∗,    Yuan Wang2,3∗,    Kairan Dou4,5,    Kaichen Zhou4 Zhanwei Zhang2,3 ,    Shuo Yang6,    Fei Tang2,3,    Jun Yin1,    Pengyu Zeng1,    Zhenzhe Ying3

### v1

@@ 4.3 Baselines
To evaluate our model, we compare it against several established baselines: (1) Foundational Model: All of our models use the Qwen-2.5-7B-Instruct model (Yang et al., 2024) as the foundational model. This includes Direct Inference, which generates answers without using retrieved context, and RAG Workflow, which guides the foundational model for RAG retrieval solely by modifying instructions, without any additional training or fine-tuning. (2) Chain-of-Thought (CoT) Methods: This category includes ReCAT and IRCoT (Trivedi et al., 2023), which enhance reasoning by explicitly generating intermediate Chain-of-Thought reasoning steps. (3) Prompt-Based Agentic RAG: This group includes models such as Self-Ask (Press et al., 2022), Iter-RetGen (Shao et al., 2023), Self-RAG (Asai et al., 2024), and Search-o1 (Li et al., 2025b), which combine retrieval and reasoning through designed prompts. We also include CR-Planner (Li et al., 2024), which employs Monte Carlo Tree Search (MCTS) for planning. (4) RL-Based Agentic RAG: This category includes models like ReARTeR (Sun et al., 2025b), R1-Search (Song et al., 2025a), and Search-R1 (Jin et al., 2025), which extend the traditional RAG paradigm by incorporating agentic search and policy learning through reinforcement learning, enabling the model to adaptively refine its retrieval and reasoning strategies during training.
@@ 4.4 Implementation Details
Our experimental framework is built upon the Qwen-2.5-7B-Instruct model (Yang et al., 2024) using the Verl framework (Sheng et al., 2025). The retrieval module utilizes the 2018 Wikipedia dump (Karpukhin et al., 2020) as the knowledge corpus, with the E5 model (Wang et al., 2024) serving as the dense retriever. For model optimization, we apply loss masking to update only the tokens generated by the model. The learning rate is set to 1e-5, with a sampling temperature of 1.0. Training is performed with a batch size of 600 (distributed across 15 NVIDIA A100 Tensor Core GPUs), generating 4 rollouts per sample and limiting the maximum retrieval count to 5. In addition, one separate GPU is used to run a 144M-parameter DistilBERT (Sanh et al., 2019) model for calculating the Evidence Quality Reward.
@@ 4.5 Main Results
On benchmarks whose distributions are aligned with the training data, EviNote-RAG achieves strong results and consistently surpasses baseline models. On HotpotQA, a dataset that requires complex multi-hop reasoning, EviNote-RAG significantly outperforms both RAG and Chain-of-Thought methods. These improvements arise from the Supportive-Evidence Notes (SEN) mechanism, which filters out spurious retrievals, and the Evidence Quality Reward (EQR), which encourages the selection of answer-critical evidence. Together, these mechanisms enable the model to construct faithful reasoning chains and maintain factual consistency, thereby yielding more accurate in-domain answers.
@@
In the more challenging out-of-domain setting, where the test distribution diverges substantially from training data, EviNote-RAG also achieves substantial gains. As shown in Tab. 1, it consistently outperforms base models, Chain-of-Thought, prompt-based, and RL-based approaches. For instance, the base model, lacking carefully designed instructions, fails under noisy retrievals and makes unstable decisions. Traditional CoT methods, such as ReCAT and IRCoT, decompose problems into reasoning steps but often rely on incomplete or irrelevant evidence, yielding weak performance on datasets like PopQA. Similarly, prompt-based approaches such as Self-Ask exhibit limited flexibility and struggle to capture critical details, as seen on 2Wiki. RL-based methods perform better, but still face difficulties with retrieval noise. For example, although Search-R1 achieves an F1 score of 0.498 on PopQA, it performs poorly on 2Wiki due to irrelevant evidence contaminating the reasoning process. By contrast, EviNote-RAG, through SEN-based filtering and EQR-guided evidence selection, maintains robustness across tasks. On TriviaQA, for example, it improves markedly over Search-R1, boosting F1 from 0.754 to 0.794. These results underscore its superior ability to generalize beyond the training distribution.
Overall, EviNote-RAG delivers consistent improvements across both in-domain and out-of-domain settings. These results demonstrate that effective noise filtering, coupled with reward design, enhances not only stability and accuracy but also generalization across diverse QA tasks.
@@ 4.6 Ablation Study
Table 2: Ablation results on in-domain and out-of-domain QA benchmarks. The table reports the contribution of Supportive-Evidence Notes (SEN) and the Evidence Quality Reward (EQR). Bold highlights the best performance, while underline marks the second best.
@@
We conducted an ablation study (Tab.2) to evaluate the impact of different model components under three configurations: Base, SEN, and SEN+EQR:
  • The Base model is Search-R1(Jin et al., 2025), an end-to-end RAG pipeline trained with reinforcement learning, encompassing the planning, search, and conclusion stages. Our model builds upon and extends this framework.
  • The Supportive-Evidence Notes (SEN) configuration modifies the prompting instructions and applies reinforcement learning constraints to enforce the required behaviors (e.g., explicitly producing a summary).
  • The SEN+EQR configuration further incorporates the Evidence Quality Reward (EQR) module, which leverages an Entail Judge model to evaluate the quality of SENs, thereby enhancing reasoning accuracy.
Across seven QA benchmarks, adding SEN to the Base system consistently improves both EM and F1, with the largest gains in out-of-domain and multi-hop settings, and steady improvements on fact-based tasks. Building on this, introducing EQR further enhances performance, particularly when reasoning chains are long or evidence is noisy. For simple factual questions, the stricter evidence filtering makes SEN and SEN+EQR comparable, but the overall trend remains clear: SEN+EQR >> SEN >> Base.
Mechanistically, these trends align with the intended roles of the modules. SEN enforces a retrieve–note–answer discipline that filters noisy retrievals, compels concise evidence summaries, and reduces variance from distractors, thereby improving the signal before generation. EQR adds entailment-based feedback that rewards notes that truly justify the final answer, curbing shallow pattern matching and guiding the policy toward faithful, stepwise reasoning. Together, they increase the salience of evidence, shorten unproductive reasoning paths, and stabilize optimization in noisy RAG environments. This shows that our training strategy effectively helps the model identify and prioritize important evidence.
In conclusion, the combination of SEN and EQR significantly enhances model performance. SEN sharpens the focus on critical evidence, while EQR improves evidence quality, leading to more accurate reasoning. Supplementary experiments (Appendix A,  B and  C) further demonstrate the superiority of our model over direct summarization and provide more detailed analyses of its effectiveness.
@@ 4.7 Training Stability and Performance Enhancement Analysis
Figure 3: Training dynamics illustrating (a) reward, (b) KL Loss, (c) Response Length, and (d) Total Times Per Token. (e) Ablation study on EQR experiments.
Stable Training Requires Proper Instructional Design.
Fig. 3(a)–(d) demonstrates the significant impact of instructional design on training stability. In our experiments, the base model (Jin et al., 2025) suffers from instability, as shown by the sharp increase in KL divergence and rapid collapse of the reward around epoch 80, as well as a sudden increase in actor loss. In contrast, EviNote-RAG, which introduces a structured retrieval-note-answer pipeline, exhibits much more stable learning curves across all metrics. Unlike traditional models that rely directly on noisy retrievals, EviNote-RAG first generates Supportive-Evidence Notes (SENs) to filter out irrelevant information and concentrate on answer-critical evidence. The integration of SEN not only improves factual consistency but also increases the probability of producing correct answers, which makes the reward denser and more reliable. Moreover, the structured instruction implicitly guides the model to adopt a reasoning process more aligned with human problem-solving habits (e.g., note-taking before answering). Such guidance reduces task difficulty, prevents the model from drifting into ill-formed reasoning or generating degenerate outputs, and stabilizes policy learning. Together, these mechanisms ensure that EviNote-RAG sustains a steady increase in reward, efficient response generation, and robust training dynamics.
Finding 1. Stable training dynamics are determined by effective instructional design, not merely by reward shaping. By generating supportive evidence notes and ensuring denser, more reliable rewards, a structured instruction like that in EviNote-RAG provides both guidance and robustness, leading to more stable training in noisy RAG environments.
Retrieval Noise Filtering Enhance Model Efficiency.
Fig. 3(c)–(d) demonstrates that EviNote-RAG achieves higher efficiency compared to the base model across multiple dimensions. First, supportive evidence note-taking filters irrelevant content early, leading to shorter and more focused responses (Fig. 3(c)). This prevents redundant reasoning and improves output relevance. Second, EviNote-RAG maintains consistently lower token-level latency (TPT), even as training progresses (Fig. 3(c)–(d)), while the base model suffers from unstable decoding behaviors such as reasoning loops and collapse around epoch 80. These improvements arise from EviNote-RAG’s structured reasoning pipeline and stabilized decoding policies, which simplified token generation and avoided unnecessary computation.
Finding 2. Retrieval noise filtering improves computational efficiency. By reducing redundant reasoning and avoiding unstable decoding loops, models achieve more concise, relevant outputs with lower latency and faster inference.
@@
Behavioral Supervision for Noise Filtering Enhances Answer Accuracy.
Tab. 2 and Fig. 3(c)–(d) further illustrate the effect of adding Evidence Quality Reward (EQR). After step 80, models with EQR continue to improve efficiency, as shown by reduced token generation time, while models without EQR exhibit longer per-token times and degraded performance (Fig. 3(c)). Moreover, EQR mitigates the risk of responses becoming excessively short, thereby avoiding degradation and the associated rise in TPT, a phenomenon evident without EQR around step 70. These results indicate that EQR enhances both efficiency and stability without sacrificing output quality. Together with supportive evidence note-taking, this highlights the broader effectiveness of behavioral supervision in refining model reasoning.
Finding 3. Supervised filtering methods, such as EQR, can improve both efficiency and output quality by preventing degenerate response patterns in RAG systems.
@@ 4.8 Case Study
To further understand the effectiveness of our method, we present a case study on the query “who wrote knock knock knocking on heaven’s door?” (Fig. 4).
Baseline: The baseline model is misled by the prominence of cover-related information in Doc 2, which frames the song as a Guns N’ Roses work. Although Doc 1 and Doc 3 mention Bob Dylan as the writer, the inconsistent framing causes the model to mistakenly output “Guns N’ Roses”.
EviNote-RAG: In contrast, our method leverages fine-grained evidence filtering and summary generation. By explicitly marking and aggregating reliable signals (e.g., multiple mentions that Bob Dylan wrote the song for the film Pat Garrett and Billy the Kid), the model successfully ignores spurious cover-related signals and outputs the correct author “Bob Dylan”.
This case study demonstrates that EviNote-RAG effectively eliminates the interference of misleading information, producing more accurate and faithful answers. More case studies are provided in Appendix D.
@@ 5 Conclusion
We present EviNote-RAG, a framework that introduces a note-taking step to extract answer-centric evidence before generation. By training LLMs to produce Supportive-Evidence Notes (SENs) and guiding them via a tailored entailment-based reward, our approach improves answer accuracy. Extensive experiments demonstrate that EviNote-RAG achieves state-of-the-art performance while enhancing training stability. These results highlight the benefits of evidence-focused abstraction for robust, faithful retrieval-augmented reasoning.
@@ A.2 Main Results
Finding 4. Effective summary strategies are not about requiring or forcing summaries, but about organizing supportive evidence to meaningfully guide reasoning.
@@
Finding 5. Gains come from highlighting key evidence. Methods that structure evidence (selection, organization, explicit marking) outperform those that only require a summary format, yielding stronger factual accuracy.
@@ B.1 Experimental Setup for Reward Shaping
Finding 6. Reward shaping is most effective when semantically aligned with evidence quality and paired with structured supervision: EQR improves over FS/FS+SR, but SEN+EQR delivers the strongest and most consistent gains across datasets.
@@ C.3 Training Stability and Efficiency
Finding 7. Semantic reward shaping alone (EQR) yields more stable training and improves factual reliability compared to stochastic shaping, but its benefits are limited without structural guidance. When combined with SEN, the two components act synergistically, producing the most consistent gains in both accuracy and generalization across QA benchmarks.

### v3

@@ 4.3 Baselines
To evaluate our model, we compare it against several established baselines: (1) Foundational Model: All of our models use the Qwen-2.5-7B-Instruct model (Yang et al., 2024) as the foundational model. This includes Direct Inference, which generates answers without using retrieved context, and RAG Workflow, which guides the foundational model for RAG retrieval solely by modifying instructions, without any additional training or fine-tuning. (2) Chain-of-Thought (CoT) Methods: This category includes ReCAT and IRCoT (Trivedi et al., 2023), which enhance reasoning by explicitly generating intermediate Chain-of-Thought reasoning steps. (3) Prompt-Based Agentic RAG: This group includes models such as Self-Ask (Press et al., 2022), Iter-RetGen (Shao et al., 2023), Self-RAG (Asai et al., 2024), and Search-o1 (Li et al., 2025b), which combine retrieval and reasoning through designed prompts. We also include CR-Planner (Li et al., 2024b), which employs Monte Carlo Tree Search (MCTS) for planning. (4) RL-Based Agentic RAG: This category includes models like ReARTeR (Sun et al., 2025b), R1-Search (Song et al., 2025a), and Search-R1 (Jin et al., 2025), which extend the traditional RAG paradigm by incorporating agentic search and policy learning through reinforcement learning, enabling the model to adaptively refine its retrieval and reasoning strategies during training.
@@ 4.4 Implementation Details
Our experimental framework is built upon the Qwen-2.5-7B-Instruct model (Yang et al., 2024) using the Verl framework (Sheng et al., 2025). The retrieval module utilizes the 2018 Wikipedia dump (Karpukhin et al., 2020) as the knowledge corpus, with the E5 model (Wang et al., 2024a) serving as the dense retriever. For model optimization, we apply loss masking to update only the tokens generated by the model. The learning rate is set to 1e-5, with a sampling temperature of 1.0. Training is performed with a batch size of 600 (distributed across 15 NVIDIA A100 Tensor Core GPUs), generating 4 rollouts per sample and limiting the maximum retrieval count to 5. In addition, one separate GPU is used to run a 144M-parameter DistilBERT (Sanh et al., 2019) model for calculating the Evidence Quality Reward.
@@ 4.5 Main Results
On benchmarks whose distributions are aligned with the training data, EviNote-RAG achieves strong results and consistently surpasses baseline models. On HotpotQA, a dataset that requires complex multi-hop reasoning, EviNote-RAG significantly outperforms both RAG and Chain-of-Thought (CoT) methods. These improvements arise from the Supportive-Evidence Notes (SEN) mechanism, which filters out spurious retrievals, and the Evidence Quality Reward (EQR), which encourages the selection of answer-critical evidence. Together, these mechanisms enable the model to construct faithful reasoning chains and maintain factual consistency, thereby yielding more accurate in-domain answers.
@@
Across out-of-domain benchmarks, EviNote-RAG achieves clear gains over the strong RL baseline (Search-R1): +91% F1 on 2Wiki (0.536 vs. 0.280, +0.256), +40% on Bamboogle (0.528 vs. 0.377, +0.151), +23% on Musique (0.336 vs. 0.274, +0.062), and +5.4% on TriviaQA (0.795 vs. 0.754, +0.040), with near-parity on PopQA (0.491 vs. 0.498). These gains are driven by behavior shaping: Supportive-Evidence Notes (SEN) compress question-conditioned evidence before generation, and the entailment-based Evidence Quality Reward (EQR) enforces that notes logically support the final answer, reducing distractor-induced errors.
Overall, EviNote-RAG delivers consistent improvements across both in-domain and out-of-domain settings. These results demonstrate that effective noise filtering, coupled with reward design, enhances not only stability and accuracy but also generalization across diverse QA tasks. The Ablations in the following section further confirm that SEN+EQR is the strongest configuration across OOD sets while controlling sequence length and token usage, aligning with our generalization claim.
@@ 4.6 Ablation Study
Settings.
Our experiments build on Base model Search-R1 (Jin et al., 2025), an end-to-end RAG pipeline trained with reinforcement learning. On top of this baseline, we examine several configurations. In the Force Summary (FS) setting, the model must output an explicit <summary> after each retrieval, with zero reward assigned if the tag is absent, which ensures strict compliance but increases reward sparsity. Relaxing this constraint, the Naive Summary (NS) setting allows the model to generate a concise <summary> of retrieved documents before answering, and the model still receives reward for a correct answer even when a summary is not provided. Building on NS, the Supportive-Evidence Notes (SEN) configuration enriches the summaries with evidence-aware annotations, improving focus and reducing misleading reasoning. Finally, the SEN+EQR configuration extends SEN by introducing the Evidence Quality Reward (EQR), which uses an Entail Judge to assess the quality of SENs and further enhance reasoning accuracy.
Overall, as shown in Tab. 2, the effectiveness of our workflow and training strategy makes SEN and SEN+EQR highly competitive, with performance consistently ranked as follows: SEN+EQR >> SEN >> Naive Summary >> Base >> Force Summary. Additionally, we observed the following experimental observations:
Effectiveness of Dynamic Summarization.
Force Summary yields inferior results, indicating that rigid structural constraints and reward sparsity hinder model adaptability. In contrast, Naive Summary significantly outperforms the baseline, demonstrating that flexible summarization improves reasoning quality. In addition, we find that changing prompts to require summarization or evidence selection does not affect model performance (see Appendix A).
Effectiveness of Evidence-aware Annotations:
The improvement observed when transitioning from Naive Summary to SEN validates our central hypothesis: the structured organization of evidence, coupled with explicit uncertainty markings (key info, uncertain info), serves as an efficient filter for noise and enhances multi-hop reasoning accuracy. SEN’s effectiveness arises from its dual capacity for selective evidence highlighting and uncertainty quantification.
Effectiveness of EQR
SEN+EQR achieves optimal performance through entailment-based supervision. The Evidence Quality Reward ensures logical consistency between generated notes and final answers, providing crucial semantic alignment that complements SEN’s structural guidance. We further demonstrate the advantages of this setting in the supplementary material, showing that it guides the model toward thorough and effective summarization behavior while maintaining stability (Appendix A,  B, and  C).
Table 2: Ablation results on in-domain and out-of-domain QA benchmarks. Bold highlights the best performance, while underline marks the second best.
@@ 4.7 Training Stability and Performance Enhancement Analysis
Figure 3: Training dynamics illustrating (a) reward, (b) KL Loss, (c) Response Length, and (d) Total Time Per Token (TPT). (e) Ablation study on EQR experiments.
Stable Training Requires Proper Workflow Design.
Fig. 3(a)–(d) highlights the crucial role of workflow design in training stability. The base model (Jin et al., 2025) collapses around epoch 80, marked by rising KL divergence, declining rewards, and unstable actor loss. In contrast, EviNote-RAG adopts a retrieval–note–answer workflow that yields consistently stable curves across all metrics. By introducing structured instructions that resemble human note-taking, it reduces task difficulty, avoids degenerate outputs, and enables more stable policy learning.
Behavioral Supervision and Noise Filtering Improve Efficiency.
Fig. 3(c) shows that incorporating evidence-quality rewards (EQR) prevents unstable performance and excessive per-token latency. This improvement arises from EviNote-RAG’s ability to produce shorter, more focused responses by filtering irrelevant content early, thereby reducing redundant reasoning. Fig. 3(d) further demonstrates that it maintains consistently lower token-level latency (Time Per Token, TPT), while the base model suffers from reasoning loops and collapses near epoch 80. Overall, the joint use of evidence note-taking and behavioral supervision yields shorter responses, reduces unnecessary computation, and enhances stability.
@@
EQR Improves SEN Quality Throughout Training.
Fig. 3(a) shows that the Evidence Quality Reward (EQR) steadily increases as training proceeds, indicating that the model learns to generate higher-quality Supporting Evidence Notes (SEN). This dynamic growth reflects the model’s improved ability to align evidence with the target answer. In addition, Fig. 3(e) demonstrates that incorporating EQR leads to SEN with stronger entailment support compared to the ablated variant, highlighting the effectiveness of behavioral supervision. Together, these results confirm that EQR not only stabilizes training but also directly enhances the reasoning quality of SEN.
@@ 4.8 Case Study
Baseline Fails by Mixing Speculation with Evidence.
In the case in Fig. 4, the baseline retrieves passages noting that Knockin’ on Heaven’s Door was performed by Guns N’ Roses but fails to distinguish between performance and authorship. Its reasoning chain introduces speculative guesses (e.g., “may suggest”), which dilute the role of explicit evidence in Doc 1 and Doc 3 stating that Bob Dylan wrote the song. As a result, the model wastes tokens on noisy deliberations and incorrectly concludes that Guns N’ Roses are the authors.
EviNote-RAG Clarify Evidence and Improve Efficiency.
EviNote-RAG highlights decisive authorship evidence (e.g., “Bob Dylan”) in key-info notes (*), keeping reasoning constrained to facts directly relevant to “who wrote …”. Moreover, the Evidence Quality Reward (EQR) guides the model to produce clearer, entailment-supported notes that isolate answer-supportive information before generation; this yields shorter answers and lower token-level latency with stable decoding. Overall gains stem from evidence shaping (SEN+EQR). More case studies are provided in Appendix D.
@@ 5 Conclusion
We present EviNote-RAG, a framework that introduces a note-taking step to extract answer-supportive evidence before answering. By training LLMs to produce Supportive-Evidence Notes (SENs) and guiding them via a tailored entailment-based reward, our approach improves answer accuracy. Extensive experiments demonstrate that EviNote-RAG achieves state-of-the-art performance while enhancing training stability. These results highlight the benefits of evidence-focused abstraction for robust, faithful retrieval-augmented reasoning. Beyond empirical gains, our work establishes a general recipe for integrating structured note-taking with reward design, offering a principled path toward more interpretable and controllable RAG systems.
@@ A.2 Main Results
Finding 1. Effective summary strategies are not about requiring or forcing summaries, but about organizing supportive evidence to meaningfully guide reasoning.
@@
Finding 2. Gains come from highlighting key evidence. Methods that structure evidence (selection, organization, explicit marking) outperform those that only require a summary format, yielding stronger factual accuracy.
@@ B.1 Experimental Setup for Reward Shaping
Finding 3. Reward shaping is most effective when semantically aligned with evidence quality and paired with structured supervision: EQR improves over FS/FS+SR, but SEN+EQR delivers the strongest and most consistent gains across datasets.
@@ C.3 Training Stability and Efficiency
Finding 4. Semantic reward shaping alone (EQR) yields more stable training and improves factual reliability compared to stochastic shaping, but its benefits are limited without structural guidance. When combined with SEN, the two components act synergistically, producing the most consistent gains in both accuracy and generalization across QA benchmarks.